# Highlights

### v2

@@ 1 Introduction
Consistency of MLEs in HMMs with finite state and observation spaces has been addressed in foundational work by Baum and Petrie (1966) and Petrie (1969). More recent work has relaxed several of the original assumptions, e.g., (Leroux, 1992; Douc et al., 2004, 2011; Genon-Catalot and Laredo, 2006; Le Gland and Mevel, 2000a, b), but these results assume time-homogeneous Markov chains. Central limit theorems for MLEs in HMMs have also been developed (Bickel et al., 1998; Douc et al., 2004; Jensen and Petersen, 1999; Brouste and Kleptsyna, 2010). Overviews of the statistical theory of HMMs can be found in Ephraim and Merhav (2002); Cappé et al. (2005). However, the specific case of a time-inhomogeneous Markov chain that maintains the same stationary distribution for the hidden chain across all time points—as occurs in the Linkage Model when $q$ is the initial distribution—has not yet been studied.
@@ 2 Models
We first define the Linkage Model for one individual and bi-allelic markers in Definition 2.1. The number of alleles on chromosome $c\in\{1,...,C\}$ at marker $m\in\{1,...,M_{c}\}$ is called $X_{c,m}\in\{0,1\}.$ We denote the genetic distance in centi Morgan (cM) between the loci m1m-1 and $m$ on chromosome $c$ by $d_{c,m}$. Recall that one centi Morgan corresponds to a $0.01$ crossing over between two genetic markers per Meiosis. Let $\mathbb{S}^{K}$ be the (K1)(K-1)-dimensional simplex and let $q:=(q_{1},\ldots,q_{K})\in\mathbb{S}^{K}$ be the ancestries of the individual from population 1,,K.1,...,K. The frequency of an allele in population $k\in\{1,...,K\}$ at marker $m$, on chromosome $c$ is called $p_{c,k,m}.$ The random variable $Z_{c,m},m=1,...,M_{c},c=1,...,C$ names the ancestral population of the allele at marker $m$ on chromosome $c$. We write $(q^{0},r^{0})$ for the true values of $(q,r)$ and $\mathbb{E},\mathbb{P}$ for the expected value and the probability respectively, with respect to the true parameters $q^{0},r^{0}.$
@@
$\mathbb{P}_{q,r}(Z_{c,m}=\tilde{k}|Z_{c,m-1}=k)$ $=\begin{cases}e^{-d_{c,m}r}+\left(1-e^{-d_{c,m}r}\right)q_{k},\textup{ if }k=\tilde{k}\\ \left(1-e^{-d_{c,m}r}\right)q_{k},\textup{ else},\end{cases}$ (2)
@@
$\mathbb{P}_{q,r}(X_{c,m}=x|Z_{c,m})=\mathbb{P}_{q,r}\left(\textup{Ber}\left(q_{Z_{c,m}}p_{c,Z_{c,m},m}\right)=x\right).$
@@
We additionally define the parameter space by $\Theta:=\{(q,r)\in[\kappa_{q},\kappa_{q}^{\prime}]^{K}\times[r_{lb},\infty]:\sum_{k=1}^{K-1}q_{k}=1\}\subset\mathbb{R}^{K+1}$.
@@
$\mathbb{P}_{q,r}\left(X^{dip}_{c,m}=x|Z^{1}_{c,m},Z^{2}_{c,m}\right)$
$\hskip 28.45274pt=\begin{cases}q_{Z^{1}_{m}}p_{c,Z^{1}_{m},m}\cdot q_{Z^{2}_{m}}p_{c,Z^{2}_{m},m},&x=2\\ q_{Z^{1}_{m}}p_{c,Z^{1}_{m},m}(1-q_{Z^{2}_{m}}p_{c,Z^{2}_{m},m})+q_{Z^{2}_{m}}p_{c,Z^{2}_{m},m}(1-q_{Z^{1}_{m}}p_{c,Z^{1}_{m},m}),&x=1\\ (1-q_{Z^{1}_{m}}p_{c,Z^{1}_{m},m})(1-q_{Z^{2}_{m}}p_{c,Z^{2}_{m},m}),&x=0.\end{cases}$
@@
$\Lambda:=-2\ln\left(\frac{\max\left\{\ell((X_{1,1},...,X_{C,M}),(q,\infty)):q\in\mathbb{S}^{K}\right\}}{\max\{\ell((X_{1,1},...,X_{C,M}),(q,r)):(q,r)\in\Theta\}}\right).$
@@ 3 Theoretical Main Results
    there exists a constant $\kappa_{d}>0$, such that $d_{c,m}\geq\kappa_{d}$ for all $m\in\{1,...,M_{c}\},c=1,...,C.$
@@
Unique MLEAssumption 3.1ConsistencyAssumption 3.1CLTAssumption 3.1,invertible Fischer Information Asymptotic Distributionof the Test StatisticAssumption 3.1
@@
Let Assumption 3.1 holdand let the Fischer information
@@
$\sqrt{M_{total}}\left((\hat{Q}^{C,M},\hat{R}^{C,M})-(q^{0},r^{0})\right)\xRightarrow{M_{total}\to\infty}\mathcal{N}\left(0,J(q^{0},r^{0})^{-1}\right).$
@@
Douc (2005) has proven the invertibility of the Fischer Information under certain constraints for stationary HMMs, i.e. in our case, if the allele frequencies and the distance between the markers are identical for every marker. This also means that the markers are either all on different chromosomes or on the same chromosome.
@@
$\Lambda\xRightarrow{M_{total}\to\infty}\chi^{2}(1)$
@@ 4.1 Simulated Data
To evaluate the statistical test in terms of Type 1 and Type 2 errors, I simulated data under both the Linkage Model and the Admixture Model for different values of $M$, $r$, and the marker distances $(d_{c,m})_{c=1,\dots,C,\ m=1,\dots,M_{c}-1}$. Specifically, I simulated 100 markers on a single chromosome with $d_{1,m}=d_{1,m+1}=:d$ for all $m\in\{1,...,M-1\}$, where $d\in\{0.1,0.5,1,2,5,10\}$ and $r\in\{1,10,100\}$. The significance level was set to $0.05$.
I computed the Type 1 and Type 2 error rates by repeating the experiment 100 times for each possible combination of $r$ and $d$. The results are shown in Figure 2. Notably, the Type 1 error remains small and consistently below the $0.05$ threshold. Moreover, both the Type 1 and Type 2 errors tend to increase with increasing values of $r$ and $d$. This behavior is expected, as larger values of $r$ and $d$ imply greater similarity between the two models, making them harder to distinguish.
In summary, the simulation results indicate that the test exhibits high power while maintaining a low Type 1 error rate.
Figure 2: Evaluation of the statistical test by using simulated data for different values of $r$ and $d$.
@@ 4.2 Real Data
I also applied the statistical test to the data from The 1000 Genomes Project Consortium (2015) with the ancestry informative marker (AIM) set by Kidd et al. (2014) consisting of $M_{total}=55$ markers on 20 different autosomes ($C=20$). For the analysis, individuals from five populations were included: Africa (AFR), Europe (EUR), South Asia (SAS), East Asia (EAS), and Admixed Americans (AMR), i.e., $K=5$. We always used all 2504 individual except the one that we are testing to calculate the allele frequencies as relative frequency. The values for $d$ for the AIM set by Kidd et al. (2014) in cM are shown in Figure 3.
Figure 3: Genetic Distances of the markers in the AIM set by Kidd et al. (2014) rounded to integer values.
The distance between loci on a single chromosome ranged from below 1 cM to approximately 129 cM.
The results of the statistical test are shown in Figure 4. In total, for 312 out of 2504 individuals (approximately 12.5%), the null hypothesis — that the Admixture Model provides a better fit to the data — cannot be rejected. Notably, there are substantial differences across populations: for individuals from Africa and East Asia, the proportion for which the null hypothesis is rejected is considerably lower compared to other groups. This highlights that it is not sufficient to define a general rule for choosing between the Linkage Model and the Admixture Model based solely on the marker set. Instead, a method — such as the statistical test defined in equation (1) — is needed, which allows model selection on an individual or population-specific basis.
Figure 4: Results of the statistical test from Definition 2.4 for the data from The 1000 Genomes Project Consortium (2015).
We further compare the covariance matrices of the MLEs obtained under the Linkage Model and the Admixture Model for individual HG00096 from Great Britain. For this analysis, I again used the AIM set introduced by Kidd et al. (2014). The results are shown in Figure 6 and Figure 6, corresponding to the Admixture and Linkage Models, respectively.
As expected, the variance of the estimators is higher in the Linkage Model due to the larger number of parameters. While both models clearly identify EUR as the primary ancestral population, the estimated admixture proportions differ significantly between models (estimated ancestry of almost 1 from Europe in the Admixture Model and estimated ancestry of approximately 0.8 from Europe in the Linkage Model).
Figure 5: Covariance Matrix for the MLE in the Admixture Model. We considered individual HG00096. The MLE for $q$ was $(1,0,0,0,0).$
Figure 6: Covariance Matrix of the MLE in the Linkage Model. We considered individual HG00096. The MLE for $q$ was $(0.8,0.04,0.04,0.08,0.04)$ and for $r$ approximately $1.8.$
@@
We also provide the code for testing whether the Linkage Model or the Admixture Model fits better to the data of $N$ individuals on GitHub. In this framework, two modeling approaches are possible: either (i) the recombination rate $r$ is assumed to vary across individuals, requiring the estimation of $N\times K$ parameters, or (ii) a single global recombination rate $r$ is assumed, reducing the number of parameters to $N\times(K-1)+1$. We chose the latter approach (ii) to reduce model complexity.
@@ 5.1 Unique MLE
$\det(B_{2})=\bigg|\begin{pmatrix}q_{1}p_{1,2}&(1-q_{1})p_{2,2}\\ 1-q_{1}p_{1,2}&1-(1-q_{1})p_{2,2}\end{pmatrix}\bigg|=q_{1}p_{1,2}-(1-q_{1})p_{2,2}.$
Hence, $B_{2}$ has the full rang if and only if $q_{1}\neq\frac{p_{1,2}}{p_{1,2}+p_{2,2}}.$ We have

### v3

# arXiv:2509.12734v3[math.ST] 27 Feb 2026

@@ 1 Introduction
Consistency of MLEs in HMMs with finite state and observation spaces has been addressed in foundational work by Baum and Petrie (1966) and Petrie (1969). More recent work has relaxed several of the original assumptions, e.g., (Leroux, 1992; Douc et al., 2004, 2011; Genon-Catalot and Laredo, 2006; Le Gland and Mevel, 2000a, b), but these results assume time-homogeneous Markov chains. Central limit theorems for MLEs in HMMs have also been developed (Bickel et al., 1998; Douc et al., 2004; Jensen and Petersen, 1999; Brouste and Kleptsyna, 2010). Overviews of the statistical theory of HMMs can be found in Ephraim and Merhav (2002); cappé2005inference. However, the specific case of a time-inhomogeneous Markov chain that maintains the same stationary distribution for the hidden chain across all time points—as occurs in the Linkage Model when $q$ is the initial distribution—has not yet been studied.
@@ 2 Models
We first define the Linkage Model for one individual and bi-allelic markers in Definition 2.1. The number of alleles on chromosome $c\in\{1,...,C\}$ at marker $m\in\{1,...,M_{c}\}$ is called $X_{c,m}\in\{0,1\}.$ We denote the genetic distance in centi Morgan (cM) between the loci m1m-1 and $m$ on chromosome $c$ by $d_{c,m}$. Recall that one centi Morgan corresponds to a $0.01$ crossing over between two genetic markers per Meiosis. Let $\mathbb{S}^{K}$ be the (K1)(K-1)-dimensional simplex and let $q:=(q_{1},\ldots,q_{K})\in\mathbb{S}^{K}$ be the ancestries of the individual from population 1,,K.1,...,K. The frequency of an allele in population $k\in\{1,...,K\}$ at marker $m$, on chromosome $c$ is called $p_{c,k,m}.$ We assume in this work that the allele frequencies are known, which is called supervised setting. The random variable $Z_{c,m},m=1,...,M_{c},c=1,...,C$ names the ancestral population of the allele at marker $m$ on chromosome $c$. We write $(q^{0},r^{0})$ for the true values of $(q,r)$ and $\mathbb{E},\mathbb{P}$ for the expected value and the probability respectively, with respect to the true parameters $q^{0},r^{0}.$
@@
$\mathbb{P}_{q,r}(Z_{c,m}=\tilde{k}|Z_{c,m-1}=k)$ $=\begin{cases}e^{-d_{c,m}r}+\left(1-e^{-d_{c,m}r}\right)q_{k},\textup{ if }k=\tilde{k}\\ \left(1-e^{-d_{c,m}r}\right)q_{\tilde{k}},\textup{ else},\end{cases}$ (2)
@@
$\mathbb{P}_{q,r}(X_{c,m}=x|Z_{c,m}=k)={p_{c,k,m}^{x}(1-p_{c,k,m})^{1-x}}.$
@@
We additionally define the parameter space by $\Theta:=\{(q,r)\in[\kappa_{q},\kappa_{q}^{\prime}]^{K}\times[r_{lb},\infty]:\sum_{k=1}^{K-1}q_{k}{\leq}1\}\subset\mathbb{R}^{K+1}$.
A visualization of the model is shown in Figure 1.
[…]
Figure 1: Visualization of the Linkage Model for two populations (blue and green) and three markers ($M=3$). (I): Reference Data Set that is used to calculate the allele frequencies. In this example we only consider bi-allelic markers. The letters A, C, G and T stand for the bases of the DNA, i.e. for Adenin, Cytosin, Guanin and Thymin. From this reference data set, we receive the allele frequencies, called $p_{k,m},k\in\{1,2\},m\in\{1,2,3\}.$ (II): Hidden Markov Model to describe the probabilities of the observed data $X_{1},X_{2},X_{3}.$ The hidden chain is represented as $Z_{1},Z_{2}$ and $Z_{3}$, the transition probabilities of state $Z_{i}$ to $Z_{i+1}$ are denoted as $A_{Z_{i},Z_{i+1}}$ and the emission probabilities $\mathbf{P}(X_{i}|Z_{i})$ are called $B_{Z_{i},X_{i}}.$ Ind. stands for Individual. (III): Estimators based on the reference data set and the Linkage Model.
@@
${\mathbb{P}_{q,r}\left(X^{dip}_{c,m}=x|Z^{1}_{c,m}=z_{1},Z^{2}_{c,m}=z_{2}\right)}$
$\hskip 28.45274pt{=}\begin{cases}{p_{c,z_{1},m}p_{c,z_{2},m},}&x=2\\ {p_{c,z_{1},m}(1-p_{c,z_{2},m})+p_{c,z_{2},m}(1-p_{c,z_{1},m}),}&x=1\\ {(1-p_{c,z_{1},m})(1-p_{c,z_{2},m})},&x=0.\end{cases}$
@@
A visualization of the Admixture Model is shown in Figure 2.
[…]
Figure 2: Visualization of the Admixture Model for two populations (blue and green) and three markers. (I): Reference Data Set that is used to calculate the allele frequencies. In this example we only consider bi-allelic markers. (II): Visualization of the model (III): Estimators based on the reference data set and the Admixture Model.
@@
$\Lambda:=-2\ln\left(\frac{\max\left\{{\mathbb{P}_{q,r}}((X_{1,1},...,X_{C,M}),(q,\infty)):q\in\mathbb{S}^{K}\right\}}{\max\{{\mathbb{P}_{q,r}}((X_{1,1},...,X_{C,M}),(q,r)):(q,r)\in\Theta\}}\right).$
@@
The test in Definition 1 compares the relative support of the Linkage Model and the Admixture Model for the observed data. Importantly, this is a model-selection test: it implicitly assumes that one of the two models is adequate. It therefore does not address the possibility that both models provide a poor description of the data, in which case selecting the better model may still yield a misleading fit. This, however, is a different but widely considered research question as e.g. considered by (Tedeschi, 2006; Mimno et al., 2015; Carstens et al., 2022).
@@ 3 Theoretical Main Results
    there exists a constant $\kappa_{d}>0$, such that $d_{m}\geq\kappa_{d}$ for all $m\in\{1,...,M\}.$
@@
Assumption 3.2 (Finitely many chromosomes).
The assumption $C=1$ is imposed solely to simplify notation. The same theory extends to the general case $C\in\mathbb{N}$. Indeed, if the number of chromosomes is finite, then the limit $M\to\infty$ necessarily implies that there are infinitely many markers on at least one chromosome. Moreover, Assumption (A5) allows only finitely many exceptions to the condition $d_{m}\leq d_{\mathrm{ub}}$. In particular, there may be up to C1C-1 indices $m$ such that $d_{m}=\infty$, which is exactly the situation corresponding to $C$ distinct chromosomes.
@@
Unique MLEAssumption 3.1ConsistencyAssumption 3.1CLTAssumption 3.1,invertible FisherInformation Asymptotic Distributionof the Test StatisticAssumption 3.1
@@
Let Assumption 3.1 hold and let the Fisher information
@@
$\sqrt{M_{total}}\left((\hat{Q}^{C,M},\hat{R}^{C,M})-(q^{0},r^{0})\right)\xRightarrow[]{M_{total}\to\infty}\mathcal{N}\left(0,J(q^{0},r^{0})^{-1}\right).$
@@
Douc (2005) has proven the invertibility of the Fisher Information under certain constraints for stationary HMMs, i.e. in our case, if the allele frequencies and the distance between the markers are identical for every marker. This also means that the markers are either all on different chromosomes or on the same chromosome. Since there is no general theory about the invertibility of the Fisher Information yet, we as Dean et al. (2014), just assume that it is invertible. For the application of the test, this is not relevant.
@@
$\Lambda\xRightarrow[]{M_{total}\to\infty}\chi^{2}(1)$
@@ 4.1 Simulated Data
To evaluate the statistical test in terms of Type 1 and Type 2 errors, I simulated data under both the Linkage Model and the Admixture Model for different values of $M$, $r$, and the marker distances $(d_{c,m})_{c=1,\dots,C,\ m=1,\dots,M_{c}-1}$. Specifically, I simulated $M=50,100,200,500,1000,10000$ markers on a single chromosome with $d_{1,m}=d_{1,m+1}=:d$ for all $m\in\{1,...,M-1\}$, where $d\in\{0.1,1,2,10\}$ and $r\in\{0.1,1,10,\infty\}$. We evaluate the performance of the test for both, diploid and haploid individuals and for $K=2,5$. The significance level was set to $0.05$.
I computed the Type 1 and Type 2 error rates by repeating the experiment 100 times for each possible combination of $r$ and $d$. The results are shown in Figure 4.
Figure 4: Evaluation of the statistical test by using simulated data for different values of $r$ and $d$ for $K=2$. The true ancestry was $q^{0}=(0.2,0.8).$ (A) $d_{m}=0.1$, (B) $d_{m}=1$, (C) $d_{m}=2$, (D) $d_{m}=10.$
Figure 4 shows that the statistical test is a level-$\alpha$ test, even for $M=50$. Moreover, for small values of $r^{0}d$, the power of the test is high and, as expected, increases with $M$. However, for large values of $r^{0}d$, the power remains low even when $M=10000$. The reason is that $\exp(-r^{0}d)\approx 10^{-5}\approx 0$ when $r^{0}d=10$. In that case, we have
$\mathbb{P}(X_{m}=x)$ $=\sum_{k=1}^{K}\mathbb{P}(X_{m}=x\mid Z_{m}=k)\mathbb{P}(Z_{m}=k)$
$\approx\sum_{k=1}^{K}p_{k,m}q_{k}=\langle p_{\cdot,m},q_{\cdot}\rangle.$
In particular, in this setting the test is unable to distinguish between the two models.
Figure 5 shows the power of the statistical test for the same cases as in Figure 4, but for diploid individuals. We see that the power is slightly higher than in the haploid case.
(B) Refer to caption
Figure 5: Evaluation of the statistical test by using simulated data for different values of $r$ and $d$ for $K=2$. We simulated diploid individuals. The true ancestry was $q^{0}=(0.2,0.8).$ (A) $d_{m}=0.1$, (B) $d_{m}=1$, (C) $d_{m}=2$, (D) $d_{m}=10.$
Figure 6 shows the performance of the statistical test for $K=5$ in the haploid setting. Again, the results are comparable to those obtained for $K=2$ in the haploid case. This suggests that the value of $K$ does not have a substantial impact on the performance of the statistical test.
Figure 6: Evaluation of the statistical test by using simulated data for different values of $r$ and $d$ for $K=5$. The true ancestry was $q^{0}=(0.2,0.05,0.25,0.4,0.1).$ (A) $d_{m}=0.1$, (B) $d_{m}=1$, (C) $d_{m}=2$, D) $d_{m}=10.$
The evaluation of the statistical test for $K=5$ and for diploid individuals are shown in Figure 15 in the Appendix.
We also used msprime Baumdicker et al. (2022) to simulate data according to an island model with $K=5$ islands. We chose the 50, 100 or 1000 markers with the hightest allele frequency difference between the populations, i.e. with the hightest
$\max\{p_{k}:k=1,...,5\}-\min\{p_{k}:k=1,...,5\}.$
With the resulting allele frequencies and distances between the markers, we then simulate the data according to the Linkage or the Admixture Model, respectively. The results are shown in Figure 8 and Figure 8. Here, the power is much higher than in the previous simulations. The reason for this is that the values of $d$ are much smaller than in the simulation directly according to the model, which is shown in Figure 11 in the appendix 7.1.2. The migration rates were chosen to achieve $F_{ST}=0.05$. The appendix 7.1.2 contains a more detailed description of the simulation with msprime and for the simulations according to the model.
Figure 7: Evaluation of the statistical test by using simulated data for different values of $r$ and $d$ for $K=5$.
Figure 8: Evaluation of the statistical test by using simulated data for different values of $r$ and $d$ for $K=5$.
@@ 4.2 Real Data
I also applied the statistical test to the data from The 1000 Genomes Project Consortium (2015) with $M=350$ loci. For the analysis, individuals from five populations were included: Africa (AFR), Europe (EUR), South Asia (SAS), East Asia (EAS), and Admixed Americans (AMR), i.e., $K=5$. To select the loci, I first extracted every 2.000 SNP and then all bi-allelic loci, where both loci have at least $5\%$ frequency in every population. This then leads to 3340 markers. I uniformly chose $350$ markers out of these 3340 markers and apply the test to these 350 markers. The distances between the markers are shown in Figure 12 in the appendix. The code for this can be found on GitHub.
The results of the statistical test are shown in Figure 9. In total, for 615 out of 2504 individuals (approximately 25%), the null hypothesis - that the Admixture Model provides a better fit to the data - cannot be rejected.
Figure 9: Results of the statistical test from Definition 2.4 for the data from The 1000 Genomes Project Consortium (2015) with $M=350$.
Interestingly, for EUR (38.77%), SAS (54%) or AMR (39.48%), the proportion of individuals for which the null hypothesis is rejected is much higher than for individuals from EAS (0.99%) or AFR (2.1%) . It might be that the linkage disequibrilium between the markers in these populations is higher in EAS and AFR or that the markers are less informative in these populations to detect linkage disequibrilium. However, these are only presumption as we cannot use the test to explain these differences between the populations. Note that we also applied the test with $M=3340$ markers. In this setting, the null hypothesis has to be rejected for all 2504 individuals.
The results in Figure 9 indicates that it is not sufficient to define a general rule for choosing between the Linkage Model and the Admixture Model based solely on the marker set. Instead, a method — such as the statistical test defined in equation (1) — is needed, which allows model selection on an individual or population-specific basis.
@@
The same method is also applicable for testing whether the Linkage Model or the Admixture Model fits better to the data of $N$ individuals. In this framework, two modeling approaches are possible: either (i) the recombination rate $r$ is assumed to vary across individuals, requiring the estimation of $N\times K$ parameters, or (ii) a single global recombination rate $r$ is assumed, reducing the number of parameters to $N\times(K-1)+1$.
Remark 4.2 (Validity of Assumption 3.1 for the true data).
We see in Figure 12 that Assumption (A4) and (A5) are met. The assumption (A2) is fulfilled for all markers and all possible combinations of the populations, while (A1) is also true due to the method how we chose the markers. However, e.g. for the $M=55$ marker in the AIM set by Kidd et al. (2014), this assumption is violated for 12 different $p_{k,m}.$ We need (A3) to justify the invertibility of the Fisher matrix, which is also true. The MLEs for $q$ was never equal to 0 or 1 and the MLE for $r$ was never equal to 0, i.e. also the assumption about the parameter space is met. Extended information about the real data set such as the precise allele frequencies and the MLEs can be found on GitHub, respectively.
Remark 4.3 (Estimation of the Uncertainty).
By comparing the two models to each other, the question which model has a lower uncertainty for the MLEs arises. This question is answered in Figure 14 and Figure 14 in the appendix by showing the covariance matrix of the MLEs. From this figures, we conclude that the uncertainty of the MLEs for $q^{0}$ is comparable small for the two models, while the uncertainty of $\hat{r}$ is relatively large.
@@ 5.1 Unique MLE
$\det(B_{2})=\bigg|\begin{pmatrix}p_{1,1}&1-p_{1,2}\\ p_{2,1}&1-p_{2,2}\end{pmatrix}\bigg|=p_{1,1}p_{2,2}-(1-p_{1,2})(1-p_{2,1}).$
Hence, $B_{2}$ has the full rang if and only if $(p_{1,1},1-p_{1,2})$ and $(p_{2,1},1-p_{2,2})$ are linearly independent from each other. We have